### v1

# Acquiring Bidirectionality via Large and Small Language Models

Abstract
Using token representation from bidirectional language models (LMs) such as BERT is still a widely used approach for token-classification tasks. Even though there exist much larger unidirectional LMs such as Llama-2, they are rarely used to replace the token representation of bidirectional LMs. In this work, we hypothesize that their lack of bidirectionality is keeping them behind. To that end, we propose to newly train a small backward LM and concatenate its representations to those of existing LM for downstream tasks. Through experiments in named entity recognition, we demonstrate that introducing backward model improves the benchmark performance more than 10 points. Furthermore, we show that the proposed method is especially effective for rare domains and in few-shot learning settings.
Acquiring Bidirectionality via Large and Small Language Models
Takumi Goto1,2, Hiroyoshi Nagao1, Yuta Koreeda1, 1Research & Development Group, Hitachi, Ltd., Tokyo, Japan, 2NARA Institute of Science and Technology, Nara, Japan, Correspondence: yuta.koreeda.pb@hitachi.com
1 Introduction
In recent years, pretrained large unidirectional language models (UniLMs), such as Llama-2 Touvron et al. (2023) and OpenAI GPT OpenAI (2024), have become readily available. Large UniLMs have demonstrated that various tasks can be solved by means of generation. On the other hand, bidirectional language models (BiLMs), most well-known by BERT Devlin et al. (2019), equipped with a classification layer are still widely used in many NLP task. In particular, BiLMs are still dominant for token-level classification tasks. For example, top three models in the CoNLL 2003 named entity recognition (NER) benchmark [^1] From Papers With Code with Wang et al. (2021); Yamada et al. (2020); Zhou and Chen (2021) being the top three as of May, 2024. Tjong Kim Sang and De Meulder (2003) and the DocRED relation extraction benchmark [^2] From Papers With Code with Ma et al. (2023); Tan et al. (2022); Xu et al. (2021) being the top three as of May, 2024. Yao et al. (2019) are all based on BiLMs.
The reason why the application of LLMs in token-level classification tasks has not progressed can be attributed to their lack of bidirectionality. As we elaborate in 2.1, the representation of a token is computed solely based on the preceding context in a UniLM. For example, in a token classification task, it is necessary to infer the label of the first word based solely on the information of that word.
In this study, we overcome this issue by proposing a method to utilize UniLMs as if they were bidirectional models. Specifically, we train a new UniLM for generating text from the end (referred to the “backward LM”) and concatenating its token representations to those of the pretrained UniLM (referred to the “forward LM”) to obtain pseudo bidirectionality. With this approach, we expect that existing LLMs can be applied not only for generation tasks but also for highly accurate solutions to token-level classification tasks.
In the experiments, we focused on named entity recognition (NER) task as an example of token-level classification tasks. We compared the NER performances with and without backward LM. As a result, we observe that adding backward LM improves the performance of the benchmarks by more than 10 points. Additionally, we demonstrated that the proposed method consistently improves performance in few-shot settings or when targeting rare domains.
The contributions of this study are as follows:
  1. 1.
    We empirically showed that unidirectionality is a problem when adopting UniLMs to token-level classification tasks.
  2. 2.
    We proposed a novel method to newly train a small-scale backward LM and concatenate its representations to those of existing LM to achieve pseudo bidirectionality in UniLMs.
2 Proposed Method
2.1 Preliminary
In this section, we review two types of LMs: UniLMs and BiLMs. Given an input sequence $\bm{x}=(x_{1},x_{2},\dots,x_{N})$ with $N$ tokens, the difference between the two models lies in how they compute representations for each word $x_{i}$ $(1\leq i\leq N)$.
In BiLMs, the representation $\mathbf{h}_{i}^{bi}$ is computed based on the context from both the beginning and the end of the text:
$\mathbf{h}_{i}^{bi}=BiLM_{\phi}(x_{i}|\bm{x}_{<i},\bm{x}_{>i}),$ (1)
where $\bm{x}_{<i}=(x_{1},x_{2},\dots,x_{i-1})$ and $\bm{x}_{>i}=(x_{i+1},x_{i+2},\dots,x_{N})$. To solve token-classification tasks, we can input $\mathbf{h}_{i}^{bi}$ to the newly added classification layers.
On the other hand, a forward LM computes the representation $\overrightarrow{\mathbf{h}}_{i}$ solely based on earlier context:
$\overrightarrow{\mathbf{h}}_{i}=\overrightarrow{Uni}LM_{\theta}(x_{i}|\bm{x}_{<i}).$ (2)
As can be seen from equation (2), UniLMs need to compute the representation for the $i$-th word without using the subsequent context $\bm{x}_{>i}$.
2.2 Utilization of the Bidirectional Language Model
The proposed method leverages the concatenated representations of both the forward LM $\overrightarrow{Uni}LM(\cdot)$ and the backward LM $\overleftarrow{Uni}LM(\cdot)$ to learn the downstream task. In contrast to the forward LM, a backward LM computes $\overleftarrow{\mathbf{h}_{i}}$ given the context from the end:
$\overleftarrow{\mathbf{h}}_{i}=\overleftarrow{Uni}LM_{\theta^{\prime}}(x_{i}|\bm{x}_{>i}).$ (3)
The final representation for the $i$-th token, considering both the forward and backward contexts, is the concatenation of $\overrightarrow{\mathbf{h}_{i}}$ and $\overleftarrow{\mathbf{h}_{i}}$, denoted as $\mathbf{h}_{i}=\mathrm{Concat}[\overrightarrow{\mathbf{h}_{i}},\overleftarrow{\mathbf{h}_{i}}]$. Therefore, the dimensions of $\mathbf{h}_{i}$ is the sum of the hidden vector dimensions of the forward and backward LMs.
To compute the concatenated representations as described above, it is necessary to share the vocabulary between the forward and backward LMs. Although there are the constraints related to the vocabulary, it is possible to use arbitrary architectures and parameter sizes for both models. For instance, we can employ a heterogeneous configurations, such as $|\theta|\gg|\theta^{\prime}|$. This means that we can utilize the existing assets of $|\theta|$ with just a small compute of training $|\theta^{\prime}|$. In this study, a part of experiments are conducted with such heterogeneous configurations.
3 Experiments
We use Llama2-7b Touvron et al. (2023) and GPT-2 (base) Radford et al. (2019) as forward LMs to assess the impact of model size. Table 1 provides an overview of the datasets used in the experiments.
3.1 Training Backward LM
Dataset # sents. Split
For training of backward LM.
Wikitext-train 1,801,350 Train
BookCorpus 74,004,228 Train
Wikitext-validation 3,760 Validation
For training of named entity recognition.
CoNLL-2003-train 14,041 Training
CoNLL-2003-valid 3,250 Validation
CoNLL-2003-test 3,453 Test
Few-NERD-train 131,767 Training
Few-NERD-valid 18,824 Validation
Few-NERD-test 37,648 Test
Table 1: Overview of the dataset used in the experiments.
Architecture Model Param. size CoNLL-2003 ($F_{1}$) Few-NERD ($F_{1}$)
BiLM BERT 124M 91.18 ± 0.17 89.50
Foward LM only GPT-2 124M 77.14 ± 0.25 88.10
      w/ Transformer Block 86.10 ± 0.63 88.71
Llama-2 7B 74.53 ± 0.32 88.05
      w/ Transformer Block 82.66 ± 0.65 87.19
Proposal GPT-2 124M + 124M 88.58 ± 0.20 89.43
Llama-2 7B + 124M 85.77 ± 0.26 89.54
Table 2: The performance on named entity recognition. For CoNLL-2003, we show the average $F_{1}$ scores on the test set, each of which is averaged over three different seeds. For Few-NERD, we report the $F_{1}$ scores on the test set, each of which is from the top performing model on the validation set amongst three different seeds due to significant standard deviation.
We train a backward LM for both of Llama2 and GPT-2, as the backward LM should have the same vocabulary as the forward LM. The architecture follows that of GPT-2 (base), but we resize the input dimension of the embedding layer to match the vocabulary size of the forward LM. We initialize the models with random parameters and train it on BookCorpus Zhu et al. (2015) and Wikitext Merity et al. (2017) (wikitext-103-raw-v1) datasets, with next token prediction objective. During the preprocessing step, we concatenate the training data from both datasets and shuffle them on a document level [^3] For Wikitext, we removed numerous empty lines and strings corresponding to headings beforehand.. Next, we perform subword tokenization with the forward LM’s tokenizer. We extract training data by cutting it into segments of 1,024 tokens, starting the beginning of the dataset, and then reversing them. For training, we set the batch size to 512 and the learning rate to 2e-5 with a cosine scheduler.
3.2 Named Entity Recognition
To evaluate the effectiveness of the proposed method in a token classification task, we focus on NER. NER aims to identify named entities such as person names, location names, and organization names in the input text. We utilize the CoNLL-2003 dataset Tjong Kim Sang and De Meulder (2003) as a general domain evaluation and Few-NERD datasets (supervised setting) Ding et al. (2021) as a rare domain evaluation. In both datasets, we adopt BIO scheme.
During training, we input the representations $\mathbf{h}_{i}$ computed by each model into the classification layer (further details will be discussed in the next section) and optimize it using cross-entropy loss. Regarding the specific training parameters, we set the batch size to 32 and employ the AdamW optimization algorithm Loshchilov and Hutter (2019). We linearly decay the learning rate starting from 1e-3. We only train the classification layer while keeping the other layers fixed. We report the score using the checkpoint that has maximum $F_{1}$ score on the validation set.
3.3 Classification Layer
The classification layer is assumed to consist of two linear layers. Let $d$ be the dimensionality of $\mathbf{h}_{i}$ and $c$ be the number of classes. We use $\mathbf{W}_{1}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{2}\in\mathbb{R}^{c\times d}$ to estimate the distribution $\mathbf{p}=\mathrm{softmax}(\mathbf{W}_{2}\mathrm{ReLU}(\mathbf{W}_{1}\mathbf{h}_{i}))\in\mathbb{R}^{c}$. As another solution to incorporate the context from the end, we also employ a fully-connected single Transformer layer as a classification layer, which is the same architecture as base type of BERT [^4] The precise formulation is described A..
3.4 Few-shot Setting
One of the potential benefits of large LMs is their ability to make generalized predictions even with a small number of training examples, leveraging the knowledge embedded in the numerous parameters. We also analyze NER performance on CoNLL-2003 with $K$-shot setting to examine the impact of limited training examples.
In our $K$-shot setting, the training data consists of $4K$ samples since we draw $K$ samples from each entity type: PER, LOC, ORG, and MISC. Note that we only extract instances from the training data that contain a single specific named entity type. During training, we set the batch size to 4 and randomly sampled the following hyperparameters; a learning rate from $\{9e-3,8e-3,\dots,2e-4,1e-4\}$ , a seed from $\{10,11,\dots,19\}$, a dropout probability from $\{0,0.1,0.2,0.3\}$. We determined top-3 hyperparameter settings in terms of $F_{1}$ score on the CoNLL-2003 validation set, and report the average $F_{1}$ on the test set of those settings.
3.5 Results
Full Dataset Setting.
We shows the results in Table 2. For BiLM setting, we train only classification layers given BERT (bert-base-cased, Devlin et al., 2019) [^5] We use the same configurations describing section 3.2 during the training.. By using the proposed method, performance has improved in both CoNLL-2003 and Few-NERD, with particularly significant improvements in scores of over 10 points in CoNLL-2003 for all models. Additionally, the settings of w/ Transformer did not outperform the proposed method. These results indicate that considering backward context improves token-classification performance and that the proposed method is better way to provide the backward context to UniLMs. In the case of Few-NERD, the results show that Llama-2 outperforms GPT-2 in the proposal setting. It can be inferred that larger forward LMs are more effective when targeting rarer domain. Finally, BERT outperforms the proposed method in most of cases, suggesting that the BiLMs are still effective for token classification tasks.
Few-shot Setting.
The experimental results using BERT(bert-base-cased) and GPT-2 (base and xl) are shown in Figure 1. The x-axis represents the number of training examples $4K$, and the y-axis represents the $F_{1}$ score on CoNLL2003 test set. The dashed lines indicate only forward LM while the solid lines represent the bidirectional and the proposal setting. As observed from Figure 1, the proposed method consistently outperforms the forward LM-only setting. Particularly, when the training data is limited, notably in less than 16 shots for each entity, the proposed method proves to be more effective than BERT performances. This has a significant value in practice as there is a valley of approaches in moderately many-shots learning; zero- to few-shots settings are more effectively addressed by in-context learning and many-shots settings are covered by BiLMs, but neither covers the middle. Addressing this valley is important, because annotating dozens of data might be justified but annotating hundreds sounds overwhelming to many practitioners.
Figure 1: Few-shot setting results on CoNLL-2003 test set. In the x-axis, the number of training examples is represented by multiplication of the number of entities ($=4$) and $K$.
4 Discussion
4.1 Effect of Model Size
Our surprising observation is that increasing the model size does not necessarily improve NER performance. Specifically, Llama2-7B lose to GPT-2 (base, 124M) in CoNLL-2003 performance (Table 2), and GPT-2 models (base 124M and xl 1.5B) have the almost same performance in few-shot experiments (Figure 1). One possible reason for this is the discrepancy in the dimensions of $\overrightarrow{\mathbf{h}_{i}}$ and $\overleftarrow{\mathbf{h}_{i}}$ when using larger forward LM. This could lead to the dominance of information from the forward LM, potentially resulting in the ineffectiveness of the proposed method. Note that the discrepancy in the dimensions cannot explain why Llama2-7B lose to GPT-2 (base) in forward LM only setting. We leave the investigation of the reason for future work.
4.2 Case Study
In the case study, it was found that the proposed method is particularly effective when there is an entity at the beginning of a sentence. Additionally, accuracy improved for the leading entity in phrases where entities are conjoined by “and.” This improvement is believed to be due to the proposed method providing context from the end. The specific examples are shown in Appendix B.
5 Related Work
A traditional approach to combining UniLMs, such as BiLSTM Schuster and Paliwal (1997) and ELMo Peters et al. (2018), is similar to the proposal method. Our study revisits this idea in the era of LLMs and aims to efficiently utilize LMs. For other approaches, Nguyen et al. (2023) formulated the loss as the sum of the losses when generating from the beginning of the text and when generating from the end of the text. Li et al. (2023b) also independently defined forward and backward models and included a normalization term to make the distributions estimated by both models closer for the same position in the text. In comparison to these studies, our method differs in its aim to utilize the latent representations of the LM and its heterogeneous composition of two types of LMs.
Li et al. (2023a) and BehnamGhader et al. (2024) fine-tunined LLMs after removing the causal attention mask to incorporate the context from the end of the sentence. Although this method requires fine-tuning for each LM, the proposed method can reuse the backward LM as long as the vocabulary and tokenizer are the same.
6 Conclusion
In this study, we proposed a method to concatenate the representations of forward and backward LM, to overcome the lack of bidirectionality problem of UniLMs. From the results in NER tasks, we could confirm the effectiveness of the proposed method. The proposed method provides more use cases to UniLM, not only for a generation model but also for an encoder model. Besides, as we discussed in Section 4.1, we need to investigate the relationship between model size and its representation quality as future work.
Acknowledgments
Input Jones Medical completes acquisition .
Foward-LM only B-PER I-ORG - - -
Proposal B-ORG I-ORG - - -
Reference B-ORG I-ORG - - -
Table 3: An examples of GPT-2 (base) in the case that an entity at the beginning of the sentence. This type of errors seems to be resolved by adding Conditional Random Field (CRF), but CRF could not resolve it.
Input Note - Lotte and Hyundai , Haitai and Samsung played two games .
Foward-LM only - - B-PER - B-ORG - - B-ORG - - - -
Proposal - - B-ORG - B-ORG - B-MISC B-ORG - - - -
Reference - - B-ORG - B-ORG - B-ORG B-ORG - - - -
Table 4: An examples of GPT-2 (base) in phrases where entities are conjoined by “and”. For the prediction corresponding to Lotte, the forward-LM struggles to infer the entity type correctly, but proposal can estimate correctly. This can be explained by the difference between considering only the context from the beginning: “Note - Lotte” or entire context, cause the improvement.
Appendices
Appendix A Classification Layer with Transformer Block
We use a one-layer Transformer layer: $Transformer(\cdot):768\mapsto 768$, which the same as BERT-base configuration. The final distribution is calculated by $\mathbf{p}=\mathrm{softmax}(\mathbf{W}_{2}^{trans}Transformer(\mathbf{W}_{1}^{trans}\overrightarrow{\mathbf{h}}_{i}))\in\mathbb{R}^{c}$, where $\mathbf{W}_{1}^{trans}\in\mathbb{R}^{768\times d}$ and $\mathbf{W}_{2}^{trans}\in\mathbb{R}^{c\times 768}$ are dense layers to align the input/output dimensions.
Appendix B Case studies
Table 3 and 4 show the actual outputs of both forward LM-only setting and proposal setting.